OpenStudy (anonymous):

The value of a "Mint State-63" (uncirculated) 1950 Jefferson nickel minted at Denver is 4/3 the value of a similar condition 1944 nickel minted at Philadelphia. Together, the value of the two coins is $28.00. What is the value of each coin?

OpenStudy (anonymous):

maybe solve \(x+\frac{4}{3}x=28\)

OpenStudy (anonymous):

you get \(\frac{7}{3}x=28\) and so \(x=\frac{3}{7}\times 28\)

OpenStudy (anonymous):

Okay thanks.

OpenStudy (anonymous):

yw